# Cosmology with Binary Neutron Stars: Does Mass-Redshift Correlation Matter?

### v2

@@ 1 Introduction
Estimating redshift only from GWs is one of the most promising techniques rely on the fact that GWs redshift mass scales (for an alternative approach using neutron star equation of state to infer redshift, see Messenger & Read (2012)). Identifying features, such as peaks or dips, in the mass distribution at different luminosity distances can estimate redshifts from a catalog of CBCs (Chernoff & Finn, 1993; Taylor et al., 2012; Taylor & Gair, 2012; Farr et al., 2019), a method known as the spectral-siren method (Ezquiaga & Holz, 2022):
@@ 2 Redshift Evolution of Features in Mass Distribution
Second, BNS formation (or the yield) has been shown to be much less sensitive to, or even unaffected by the formation metallicity. This is in stark contrast to BBH formation, where has repeatedly been shown to drop at high metallicities. This behavior is consistently observed across various population synthesis models with a very broad variation of physics assumptions (including BPASS, MOBSE, StarTrack, COMPAS, SEVN Eldridge & Stanway, 2016; Giacobbo & Mapelli, 2018; Giacobbo et al., 2018; Klencki et al., 2018; Neijssel et al., 2019; Santoliquido et al., 2021; Broekgaarden et al., 2022; Iorio et al., 2023; Fishbach & van Son, 2023). DCO formation through the CE channel is influenced by metallicity-dependent winds, which reduce the progenitor core masses . For BBHs, lower-mass BHs receive larger natal kicks because these kicks are assumed to scale inversely with mass . Hence this leads to a loss of potential BBHs at high metallicities. In contrast, neutron star progenitors are not assumed to experience mass-scaled natal kicks, and thus BNS formation is not sensitive to metallicity in the same way (see for a more in depth discussion of the origin and robustness of this result). This lack of metallicity dependence suggests that the BNS mass distribution will be less impacted by the changing metallicity distribution of star formation across redshifts (Chruślińska, 2024).
@@ 3 Spectral Siren Cosmology with an Evolving BNS Mass Spectrum
To test whether assuming a redshift-independent BNS mass population biases our cosmological inference, we conduct an injection-recovery campaign by generating mock GW data for BNS mergers. The mock GW data is generated in Cosmic Explorer sensitivity (Evans et al., 2021), with the injected population assumed to be fiducial and Mandel-Müller. We select the Mandel-Müller model because it exhibits the highest mass-redshift correlation (see . If this model still allows for unbiased cosmological parameter inference, then the other models should also be safe.
@@ Appendix D Method
For the uncorrelated case, we model the mass population, marginalized over redshift, using a two-dimensional Gaussian Mixture Model, $f^{\text{GMM}}_{\text{2D}}(M_{c},q)$, and the redshift population, marginalized over masses, using a one-dimensional Gaussian Mixture Model, $f^{\text{GMM}}_{\text{1D}}(z)$. The uncorrelated population is then given by:

### v3

@@ 1 Introduction
Estimating redshift only from GWs is one of the most promising techniques relies on the fact that GWs redshift mass scales (for an alternative approach using neutron star equation of state to infer redshift, see Messenger & Read (2012)). Identifying features, such as peaks or dips, in the mass distribution at different luminosity distances can estimate redshifts from a catalog of CBCs (Chernoff & Finn, 1993; Taylor et al., 2012; Taylor & Gair, 2012; Farr et al., 2019), a method known as the spectral-siren method (Ezquiaga & Holz, 2022):
@@ 2 Redshift Evolution of Features in Mass Distribution
Second, BNS formation (or the yield ) has been shown to be much less sensitive to, or even unaffected by the formation metallicity. This is in stark contrast to BBH formation, where has repeatedly been shown to drop at high metallicities. This behavior is consistently observed across various population synthesis models with a very broad variation of physics assumptions (including BPASS, MOBSE, StarTrack, COMPAS, SEVN Eldridge & Stanway, 2016; Giacobbo & Mapelli, 2018; Giacobbo et al., 2018; Klencki et al., 2018; Neijssel et al., 2019; Santoliquido et al., 2021; Broekgaarden et al., 2022; Iorio et al., 2023; Fishbach & van Son, 2023). Double compact object formation through the common envelope channel is influenced by metallicity-dependent winds, which reduce the progenitor core masses . For BBHs, lower-mass BHs receive larger natal kicks because these kicks are assumed to scale inversely with mass . Hence this leads to a loss of potential BBHs at high metallicities.
@@
We observe that the correlations between mass and redshift across these models are weak, supporting our assumption that the mass distribution is nearly independent of redshift. Among all the variations, the Mandel & Müller (2020) mass distribution exhibits the strongest variation with redshift.
@@ 3 Spectral Siren Cosmology with an Evolving BNS Mass Spectrum
whether assuming a redshift-independent BNS mass population biases our cosmological inference; we conduct an injection-recovery campaign by generating mock GW data for BNS mergers. The mock GW data is generated in Cosmic Explorer sensitivity (Evans et al., 2021), with the injected population assumed to be fiducial and Mandel-Müller. We select the Mandel-Müller model because it exhibits the highest mass-redshift correlation (see . If this model still allows for unbiased cosmological parameter inference, then the other models should also be safe.
@@ Appendix D Method
For the uncorrelated case in Fig. 5, we model the mass population, marginalized over redshift, using a two-dimensional Gaussian Mixture Model, $f^{\text{GMM}}_{\text{2D}}(M_{c},q)$, and the redshift population, marginalized over masses, using a one-dimensional Gaussian Mixture Model, $f^{\text{GMM}}_{\text{1D}}(z)$. The uncorrelated population is then given by: